Together, a computer, monitor, and printer cost $1,935. The monitor costs $400 less than the computer, and the printer costs $50 more than the monitor. Find the cost of the printer.
Answer by drj(1380)

Together, a computer, monitor, and printer cost $1,935. The monitor costs $400 less than the computer, and the printer costs $50 more than the monitor. Find the cost of the printer.

Step 1. Let x be the cost of the monitor, let x+50 be the cost of the printer, x+400 be the cost of the computer.

Step 2. x+x+50+x+400=1935 Total cost of computer, monitor, and printer. Then

3x-450=1935

3x=1935-450

3x=1485

x=495

With x=495, then x+50=545 and x+400=895. Check to see if total is 1935 or 495+545+895=1935 which is a true statement.

Step 3. ANSWER. The cost of the printer is $545.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J